# Effect of Overheat and Direct Flow Loading on the MEMS Bistable Flow Sensor

### v1

@@ Abstract
We present the findings from an experimental study of a MEMS flow sensor, in which an initially curved, double-clamped bistable microbeam serves as the primary sensing element. Our research focuses on exploring how the overheat ratio, direct flow loading, and turbulence-induced vibration, affect the sequential snap-through (ST) buckling and snap-back (SB) release of an electrostatically actuated beam heated by an electric current. The sensor is fabricated from highly doped single-crystal silicon using a silicon-on-insulator (SOI) wafer. Positioned at the chip’s edge, the microbeam is exposed to airflow, enabling concurrent dynamic response measurements, with a laser Doppler vibrometer and a video camera.
Our research demonstrates that for this sensing principle, the overheat ratio can be significantly lower compared to conventional thermal sensing elements, pointing to the potential for substantial energy savings. We also emphasize the significant impact of flow angles and vibrations on the critical ST and SB voltages, which are vital for the flow sensor’s output. Additionally, we introduce the previously inaccessible experimental observation of the beam profile’s time history during the snap-through transition.
The potential impact of this research lies in the development of more robust MEMS flow sensors with enhanced sensitivity and a better understanding of their response to environmental factors, which could have broader applications in fields such as aerospace, environmental monitoring and industrial process control.
@@ 1 Introduction
Microscale flow sensors play a key role in advancing scientific and technological endeavors. Among the various types of flow sensors, those based on microelectromechanical systems (MEMS) have become increasingly important for monitoring gas flow in industrial environments, sensing environmental flows, detecting marine conditions and analyzing biomedical flows ho1998micro . These sensors have a wide range of applications in technical fields, such as “smart skin”, the Internet of Things (IoT), and flight control (e.g., dong2017mems ; shi2020progress ).
@@
The utilization of MEMS-based flow sensors presents several advantages, including compactness, power efficiency, high sensitivity, and compatibility with mass production requirements. Consequently, a multitude of MEMS flow sensing methods have emerged over the past three decades wang2009mems ; kuo2012micromachined ; ejeian2019design . These sensors operate based on either thermal or mechanical principles, with a majority of the previously developed flow sensors relying on the thermal principle ejeian2019design . Thermal flow sensors offer robust and straightforward manufacturing processes, linear sensor characteristics, and lack of moving parts, distinguishing them from mechanical flow sensing methods. However, their major drawbacks include high power consumption and undesired heat dissipation. Recent advancements in microelectronic device design and fabrication have aimed to mitigate these limitations of thermal flow sensors le2021fabrication . Nonetheless, mechanical flow sensors continue to play a vital role in water flow, especially for laminar flow. For example, sensors based on micro/nano-cantilevers have found diverse applications in chemical and biological sensing mathew2018review . Mechanical flow sensors often rely on drag or lift force sensing kalvesten1996integrated ; svedin1998lift ; svedin2003new , frequency monitoring bouwstra1990resonating , or artificial hair sensor designs dagamseh2011bio . Although various MEMS flow sensors are under research, specific sensing capabilities are necessary for each technical application, contingent on environmental and flow conditions.
@@
Previous studies krylov2018IEEE ; kessler2018JMEMS introduced an alternative flow sensing principle,combining thermal and mechanical responses in a bifurcation-type sensor. These studies showed that the electrostatically and electrothermally (Joule heated) actuated initially curved double-clamped silicon microbeam exhibits bistable behavior, in other words, remaining in one of the two equilibrium position states under the same load. As a result of this bistability, the curved beam shows high responsiveness to external stimuli near two equilibrium states and has great potential for sensor implementation. This mechanical structure is pushed to its stability limits, and is highly sensitive to even the slightest changes in heat transfer caused by flow. As a result, reducing the overheat ratio (the temperature difference between the beam and the surrounding airflow) can reduce power consumption. In addition, a curved microbeam device at the chip’s edge can be used as a spatial velocity gradient sensor kessler2020flow .
@@
The feasibility of employing such devices for airflow sensing has been confirmed through a combination of theoretical analysis and experimental testing. However, to effectively utilize the laboratory prototype in turbulent airflow measurements within practical scenarios, it is imperative to enhance our understanding of the device’s response to various influencing factors. These factors include overheat ratios (the absolute temperature to which the microbeam is heated as a multiple of the ambient temperature), the sensitivity of the microbeam to direct flow loading, as well as noise and vibrations that may arise during turbulent flow measurements, such as those over aircraft wings or wind turbine blades.
@@
This study aims to investigate into how the sensor interacts with these factors. To achieve this, we have expanded the diagnostic system introduced in kessler2020flow to incorporate a multipoint Doppler laser vibrometer. This addition allows us to visualize the beam shape during transitions between two equilibrium states. The outcome of this extension has revealed a previously only theoretically predicted phenomenon – asymmetric snap-through motion, as described in medina2012symmetry . Furthermore, this finding sheds light on the limitations of the model presented in the paper, emphasizing the need for further refinements.
@@ 2.1 Sensor and chip design
The bistable flow sensor based on an initially curved microbeam was introduced in  kessler2020flow ; kessler2021sampling . We briefly review the device design and its operational principle for clarity and completeness. We use an initially curved bell-shaped electrostatically actuated double-clamped beam, as shown in Figure 1, as the sensing element. The beam is fabricated from the device layer of a silicon-on-insulator (SOI) wafer, using the SOIMUMPsTM process soimumps . The beam, anchors and electrode are all made of highly doped single-crystal silicon with a resistivity of 0.02 $\Omega\cdot$cm. The beam is designed to perform an in-plane, parallel to the substrate, motion and to deflect in the $z$-direction. The nominal, as designed dimensions of the beam are given in Table 1 and definitions are according to Figure 1.
Figure 1: (a) A scanning electron microscope micrograph of the device (the scale bar is 100 $\mu m$) and device 3D schematics. The double-clamped curved beam is shown in the initial upward curved configuration. The values $V_{\mathrm{h}}$ and $V_{\mathrm{ES}}$ are the heating and electrostatic voltages, respectively (G is a ground). (b) Schematic front view of the beam, its main geometric parameters, and the airflow direction (not to scale).
@@
The operational principle of the device relies on coupled thermal and mechanical sensing. It begins with an initially curved bistable microbeam, subjected to controlled heating, using a weak electric current regulated by a constant heating voltage, denoted as $V_{\mathrm{h}}$. Simultaneously, cooling is achieved through the presence of an airflow, as illustrated in Figure 1. The voltage difference $V_{\mathrm{h}}$ is applied across the two beam anchors, with one of them serving as the ground (G in Fig. 1a).
The time-periodic electrostatic force acting on the beam is controlled by the actuating voltage $V_{\mathrm{ES}}$, applied to an electrode, as marked in Figure 1. Increasing $V_{\mathrm{ES}}$ beyond a critical snap-through (ST) threshold, denoted as $V_{\mathrm{ST}}$, triggers a dynamic shift from the microbeam’s first to the second stable equilibrium position. Conversely, reducing the voltage below the lower critical threshold, $V_{\mathrm{SB}}$, causes the beam to snap back (SB). The interaction between Joule heating and cooling via the airflow influences the curvature of the beam and the critical voltage values. Measurements of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ are used to determine the air velocity. This device offers enhanced performance owing to its enhanced sensitivity near the ST and SB points krylov2018IEEE ; kessler2018JMEMS . Additionally, it possesses a unique capability to measure flow at two closely spaced points at two stable positions of the microbeam, allowing for the assessment of flow velocity gradients, as discussed in kessler2020flow .
@@ 2.2 Setup
The experimental setup is illustrated in Figure 2(a). During the experiments, an AFG3022C waveform generator from Tektronix is employed. It provides a voltage through channel 1, which is subsequently amplified through a signal amplifier (Trek PZD350A). This amplified voltage is then used to apply $V_{\mathrm{ES}}$  to the electrode. Additionally, to generate the electric current required for Joule heating within the beam, a constant heating voltage $V_{\mathrm{h}}$ is applied across the beam’s anchors, with this voltage being supplied by the same waveform generator through channel 2.
Figure 2: (a) Overview of the experimental setup. (b) Optical image of the experimental setup and (c) schematics of the chip from the perspective of the side camera, with the laser from the LDV, and the airflow at an angle $\alpha^{\circ}$ relative to the beam.
The real-time displacements of the beam’s center are recorded using a single-point laser Doppler vibrometer (LDV, Polytec, OFV534). The LDV focuses its laser beam on the side face of the beam, which has a width of $b=25\;\mu$m. The values of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ are extracted from the time series recorded with an oscilloscope (DSOX2004A, Keysight), synchronously with the LDV signal.
To regulate the flow rate, we employ a pressure gauge (P-30PSIG-D/5P, Alicat) to monitor the pressure in the air supply system. A rectangular nozzle (1” High Power, EXAIR) generates a wide, uniform jet airflow directed towards the microbeam. The air supply is calibrated using a commercial anemometer (CTV 110, Kimo). Our experiments encompass a range of air velocities spanning from 0 to approximately 15 m/s. We estimate the systematic uncertainty of these measurements to be 2.8 m/s. Additionally, it is worth noting that the airflow can be applied at various angles, denoted as $\alpha^{\circ}$ in Figure 2.
The beam is strategically positioned at the chip’s edge, as depicted in Figure 2(b,c). This positioning offers the advantage of simultaneously capturing real-time beam velocity from a top view through LDV, while also enabling direct imaging of ST and SB events using a custom-built optical arrangement. The optical setup comprises a video camera (UI-6250SE-M-GL, IDS) and a horizontally mounted tube microscope (Navitar). In most of our experiments, the side camera’s use is primarily limited to the initial setup preparations, as the bulk of quantitative measurements are conducted through the LDV.
@@ 2.3 Measurement methodology
Figure 3(a) illustrates the time history of a cyclic voltage $V_{\mathrm{ES}}$ applied to the electrode (labeled in blue, left vertical axis) and the corresponding $z$ value of the beam midpoint ($x=L/2$), which is obtained using LDV (labeled in red, right vertical axis). The electrostatic voltage signal $V_{\mathrm{ES}}$ is a linear ramp with a constant rate of about 170 V per period. The analog output from the LDV controller and the input voltage signal are looped back into a digital storage oscilloscope, allowing for the synchronized recording of their time histories. The determination of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ values occurs during the post-processing phase. This is achieved by identifying the displacement threshold transition events within the recorded data, as illustrated in Figure 3(b). These threshold events correspond to a 5 $\mu$m jump in displacement, serving as the triggers for ST and SB events.
@@
The actuation process is repeated for four cycles, operating the beam at a frequency of 400 Hz. Averaging is employed to mitigate data variability, facilitating the calculation of the mean value and its associated standard error. It is important to note that the mechanical response can be considered quasi-static due to the significantly lower frequency of the electrical signal chosen, which is of the order of $\mathcal{O}(10^{5})$ Hz compared to the fundamental frequency of the beam.
The time history of the beam center displacement is presented in Figure 3(c) for different $V_{\mathrm{h}}$ voltage levels. Crossing of dashed lines with the electrostatic voltage linear ramp signal help to indicate the values of $V_{\mathrm{ST}}$  and $V_{\mathrm{SB}}$ ,at the moments of transitions, for the case of the heating voltage $V_{\mathrm{h}}$ equal to 0 V. A closer examination of the zoom in view in Figure 3(d) reveals a noticeable shift in the time when the threshold is crossed for increasing heating voltage $V_{\mathrm{h}}$ in the range 0 $\div$ 1.66 V. This shift results in distinctly different critical electrostatic voltages $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$. We will address this shift quantitatively in the following, but we first address the bias in these measurements, due to the LDV heating of the beam.
@@ 2.4 LDV-Induced Beam Heating Bias
The specific setup explained above enables us to obtain unique results, including direct measurement of the beam’s response time history during the transition using LDV. However, the LDV system comes with a drawback known as the “loading effect,” which inadvertently heats the beam during measurements, as discussed in kassie2020effect . To assess the bias on the measured $V_{\mathrm{ST}}$  and $V_{\mathrm{SB}}$  values, we employ digital camera imaging of the beam and compare the results obtained with and without LDV. It is important to note that the image processing method for detecting ST and SB events is less precise compared to the LDV approach. Therefore, in these experiments, we extended the period to 60 seconds to achieve more accurate results. This extension was necessary because of the video camera’s lower temporal resolution, which demanded a longer measurement duration for precise observations. Consequently, the bias due to LDV heating could be larger than in the case of high-frequency measurements in our experiments, due to a significantly longer time period of the measurement.
The figures in Figure 4(a, b) show the results with “LDV on” and
Figure 4: (a) Critical snap-through voltage ($V_{\mathrm{ST}}$) and (b) critical snap-back voltage ($V_{\mathrm{SB}}$) as a function of the flow velocity $U$. Circle markers indicate measurements with LDV, while square markers indicate measurements without LDV. (c) Sensitivity of the snap-back event, a dimensionless measure of the change in voltage with respect to the reference voltage $V_{\mathrm{SB}}$ at $U=0$ m/s. Dashed lines represent a quadratic fit.
“LDV off,” represented by circles and squares. Using the LDV system increases both $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ by around 4 V and 6 V, respectively. To further quantify the influence of LDV heating, we examine the relative change in SB voltage, denoted as $\Delta V_{\mathrm{SB}}=100\times(V_{\mathrm{SB}}-V_{\mathrm{SB}}|{U=0})/V_{\mathrm{SB}}|{U=0}$ in Figure 4(c). Notably, $\Delta V_{\mathrm{SB}}$ increases from 6% to 10% at an airflow velocity of 12 m/s when the LDV is active. We will discuss this bias further in Section 3.1, where we explore how Joule heating affects the behavior of the flow velocity sensor.

### v2

@@ Abstract
We present the findings from an experimental study of a MEMS flow sensor in which an initially curved, double-clamped bistable microbeam is the primary sensing element. Our research explores how the overheat ratio, direct flow loading, and turbulence-induced vibration affect the sequential snap-through (ST) buckling and snap-back (SB) release of an electrostatically actuated beam heated by an electric current. The sensor is fabricated from highly doped single-crystal silicon using a silicon-on-insulator (SOI) wafer. Positioned at the chip’s edge, the microbeam is exposed to airflow, enabling concurrent dynamic response measurements with a laser Doppler vibrometer and a video camera.
Our research demonstrates that the overheat ratio can be significantly lower for this sensing principle than conventional thermal sensing elements, pointing to the potential for substantial energy savings. We also emphasize the significant impact of flow angles and vibrations on the critical ST and SB voltages, which are vital for the flow sensor’s output. Additionally, we introduce the first direct experimental observation of the beam profile’s time history during the snap-through /snap-back transition.
The potential impact of this research lies in developing more robust MEMS flow sensors with enhanced sensitivity and a better understanding of their response to environmental factors, which could have broader applications in fields such as aerospace, environmental monitoring, and industrial process control.
@@
{graphicalabstract}[Uncaptioned image]
{highlights}
A MEMS bistable flow sensor based on the initially curved double-clamped microbeam located at the edge of the chip is presented
Influence of overheat ratio, direct flow loading, and turbulence-induced vibrations on the critical voltages are investigated
Direct beam profile measurement during the dynamic snap-through and snap-back transitions is presented.
@@ 1 Introduction
Microscale flow sensors play a key role in advancing scientific and technological endeavors. Among the various types of flow sensors, those based on microelectromechanical systems (MEMS) have become increasingly important for monitoring gas flow in industrial environments, sensing environmental flows, detecting marine conditions, and analyzing biomedical flows [1]. These sensors have a wide range of applications in technical fields, such as “smart skin”, the Internet of Things (IoT), and flight control (e.g., [2, 3]).
@@
The utilization of MEMS-based flow sensors presents several advantages, including compactness, power efficiency, high sensitivity, and compatibility with mass production requirements. Consequently, a multitude of MEMS flow sensing methods have emerged over the past three decades [4, 5, 6]. These sensors operate based on thermal or mechanical principles, and most previously developed flow sensors are based on the thermal principle [6]. Thermal flow sensors offer robust and straightforward manufacturing processes, linear sensor characteristics, and lack of moving parts, distinguishing them from mechanical flow sensing methods. However, the main drawbacks of the thermal sensors include high power consumption, calibration requirements, unwanted heat dissipation, and limitations in miniaturizing heat sources. The reduction in the size of the sensing element was the subject of extensive research endeavors  [7]. Recent advances in the design and fabrication of microelectronic devices have focused on reducing the power consumption of thermal flow sensors [8, 9]. However, mechanical flow sensors continue to play a vital role in water flow, especially in laminar flow. For example, micro/nanocantilever-based sensors have found various applications in chemical and biological sensing [10]. Mechanical flow sensors often rely on drag or lift force sensing [11, 12, 13], frequency monitoring [14], or artificial hair sensor designs [15]. While mechanical flow sensors are sensitive enough to detect low flow rates, they face challenges in measuring higher flow rates. Although various architectures of MEMS flow sensors are being developed in both industry and academia, specific sensing capabilities are sought to meet the technical requirements of each application, depending on environmental and flow conditions.
@@
Previous studies [16, 17] introduced an alternative flow sensing principle that combines thermal and mechanical responses in a bifurcation-type sensor. These studies showed that the electrostatically and electrothermally (Joule heated) actuated initially curved double-clamped silicon microbeam exhibits bistable behavior, in other words, remaining in one of the two equilibrium position states under the same load. As a result of this bistability, the curved beam shows high responsiveness to external stimuli near two equilibrium states and has excellent potential for sensor implementation. This property of bistable systems plays an instrumental role in many MEMS/NEMS-based devices [18, 19]. When the mechanical structure is pushed to its stability limits, it becomes highly sensitive to even the slightest changes in heat transfer caused by flow. Reducing the overheating ratio (which is the ratio between the absolute temperature to which the microbeam is heated and the ambient temperature) can significantly decrease power consumption. Direct comparison of different sensors is challenging due to their completely different operating principles. However, note that the measured sensitivity of a bifurcation-type sensor based on the bistable curved silicon microbeam (1.9 V/m/s [16]) is comparable to or better than the state-of-the-art values reported for thermal flow sensors [20].
@@
The feasibility of employing such devices for airflow sensing has been confirmed through theoretical analysis and experimental testing [16, 17, 21, 22]. However, enhancing our understanding of the device’s response to various influencing factors is imperative to effectively utilize the laboratory prototype in airflow measurements within practical scenarios. These factors include overheat ratios, the sensitivity of the microbeam to direct flow loading, and noise and vibrations that may arise during turbulent flow measurements, such as those over aircraft wings or wind turbine blades.
@@
This study aims to investigate how the sensor interacts with these factors. An especially important result is the direct visualization of the beam’s shape during transitions between two equilibrium states. We added a multi-point Doppler laser vibrometer to the diagnostic system introduced in [21] to achieve this. The outcome of this extension has revealed a previously theoretically predicted phenomenon – asymmetric snap-through visualization of an electrostatically actuated beam, as described in [23]. Furthermore, this finding sheds light on the limitations of the mathematical model presented in the paper, highlighting the need for further refinements.
@@ 2.1 Sensor and chip design
The bistable flow sensor based on an initially curved microbeam was introduced in  [21, 22]. We briefly review the device design and its operational principle for clarity and completeness. We use an initially curved bell-shaped electrostatically actuated double-clamped beam, as shown in Figure 1(a), as the sensing element. The beam is fabricated from the device layer of a silicon-on-insulator (SOI) wafer using the SOIMUMPsTM process [24]. The beam, anchors, and electrode are all highly doped single crystal silicon with a resistivity of 0.02 $\Omega\cdot$cm. The beam is designed to perform an in-plane motion parallel to the substrate and to deflect in the $z$ direction. The nominal dimensions of the beam “as designed” dimensions are given in Table 1, and their definitions are shown in Figure 1(b).
Figure 1: (a) A scanning electron microscope micrograph of the device, and (b) Schematic view of the beam, its main geometric parameters, and the airflow direction (not to scale). The double-clamped curved beam is shown in the initial upward curved configuration. The values $V_{\mathrm{h}}$ and $V_{\mathrm{ES}}$ are the heating and electrostatic voltages, respectively (G is a ground).
@@
The operational principle of the device relies on coupled thermal and mechanical sensing. It begins with an initially curved bistable microbeam, subjected to controlled Joule heating, using a weak electric current regulated by a constant heating voltage, denoted as $V_{\mathrm{h}}$. The voltage difference $V_{\mathrm{h}}$ is applied across the two beam anchors, with one of them serving as the ground (G). Simultaneously, airflow removes heat from the microbeam, as illustrated in Figure 1(b).
The electrostatic time-periodic force acting on the beam is controlled by the actuating voltage $V_{\mathrm{ES}}$, applied to an electrode, as marked in Figure 1. Increasing $V_{\mathrm{ES}}$ beyond a critical snap-through (ST) voltage, denoted as $V_{\mathrm{ST}}$, triggers a dynamic shift from the microbeam’s first to the second stable equilibrium position. On the contrary, reducing the voltage below the lower critical value $V_{\mathrm{SB}}$, causes the beam to snap back (SB). The interaction between Joule heating and airflow cooling influences the beam’s curvature and the critical voltage values. Measurements of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$enable an estimate of air velocity. The enhanced performance is due to the high sensitivity of the bistable beam near the ST and SB points [16, 17]. Additionally, the device in its present design possesses a unique capability to measure the flow velocity at two closely spaced points corresponding to two stable positions of the microbeam, thus allowing for the direct measurement of the flow velocity gradients, as discussed in [21].
@@ 2.2 Setup
The experimental setup is illustrated in Figure 2(a). During the experiments, an AFG3022C waveform generator from Tektronix is employed. It provides a voltage through channel 1, subsequently amplified through a signal amplifier (Trek PZD350A). This amplified voltage is then used to apply $V_{\mathrm{ES}}$  to the electrode. Furthermore, to generate the electric current required for Joule heating within the beam, a constant heating voltage $V_{\mathrm{h}}$ is applied across the beam’s anchors, with this voltage being supplied by the same waveform generator through channel 2.
Figure 2: (a) Overview of the experimental setup. (b) Optical image of the experimental setup and (c) schematics of the chip from the perspective of the side camera, with the laser from the LDV, and the airflow at an angle of attack () relative to the beam.
The real-time displacements of the beam’s center are recorded using a single-point laser Doppler vibrometer (LDV, Polytec, OFV534). The LDV focuses its laser beam on the side face of the beam, which has a width of $b=25\;\mu$m. The values of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ are extracted from the time series recorded with an oscilloscope (DSOX2004A, Keysight) synchronously with the LDV signal.
To regulate the flow rate, we employ a pressure gauge (P-30PSIG-D/5P, Alicat) to monitor the pressure in the air supply system. A rectangular nozzle (1” High Power, EXAIR) generates a wide, uniform jet airflow toward the microbeam. The air supply is calibrated using a commercial anemometer (CTV 110, Kimo). Since the spatial resolution of the anemometer cannot ensure an accurate measurement of the uniform velocity distribution within the jet core, especially when the sensor is placed at a distance from it, we have experimentally estimated the uncertainty of the sensor position with respect to the flow nozzle and consequently the uncertainty of the velocity near the sensor. We estimate the maximum uncertainty of the velocity measurement near the sensor to be $\pm$2.8 m/s mainly to demonstrate the feasibility and working principle of the velocity sensor. Our experiments cover a range of air velocities ranging from 0 to 15 m/s. Additionally, it is worth noting that the airflow can be applied at various angles of attack, shown in Figure 2.
The beam is strategically placed at the edge of the chip, as depicted in Figure 2(b,c). This positioning offers the advantage of simultaneously capturing real-time beam velocity from a top view through LDV, while also enabling direct imaging of ST and SB events using a custom-built optical arrangement. The optical setup comprises a video camera (UI-6250SE-M-GL, IDS) and a horizontally mounted tube microscope (Navitar). In most of our experiments, the side camera’s use is primarily limited to the initial setup preparations, as the bulk of quantitative measurements are conducted through the LDV.
@@ 2.3 Measurement methodology
Figure 3(a) illustrates the time history of a cyclic voltage $V_{\mathrm{ES}}$ applied to the electrode (labeled in blue, left vertical axis) and the corresponding $z$ value of the beam midpoint ($x=L/2$), which is obtained using LDV (labeled in red, right vertical axis). The electrostatic voltage signal $V_{\mathrm{ES}}$ is a linear ramp with a constant rate of about 170 V per period. The analog output from the LDV controller and the input voltage signal are looped back into a digital storage oscilloscope, allowing for the synchronized recording of their time histories. The determination of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ values occurs during the post-processing phase. This is achieved by identifying the displacement threshold transition events within the recorded data, as illustrated in Figure 3(b). These threshold events correspond to a 5 $\mu$m jump in displacement, serving as triggers for ST and SB events.
@@
The actuation process is repeated for four cycles, operating the beam at 400 Hz. Averaging is used to mitigate the variability of the data, facilitating the calculation of the mean value and its associated standard error. It is important to note that the mechanical response can be considered quasi-static due to the significantly lower frequency of the chosen electrical signal, which is on the order of $\mathcal{O}(10^{5})$ Hz compared to the fundamental frequency of the beam.
The time history of the beam center displacement is presented in Figure 3(c) for different $V_{\mathrm{h}}$ voltage levels. Crossing of dashed lines with the electrostatic voltage linear ramp signal helps to indicate the values of $V_{\mathrm{ST}}$  and $V_{\mathrm{SB}}$ , at the moments of transitions, for the case of the heating voltage $V_{\mathrm{h}}$ equal to 0 V. A closer examination of the zoom in view in Figure 3(d) reveals a noticeable shift in the time when the threshold is crossed for increasing heating voltage $V_{\mathrm{h}}$ in the range 0$\sim$1.66 V. This shift results in distinctly different critical electrostatic voltages $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$. We will address this shift quantitatively in the following, but we first address the bias in these measurements due to the LDV heating of the beam.
@@ 2.4 LDV-Induced Beam Heating Bias
The specific setup explained above enables us to obtain unique results, including directly measuring the beam response time history during the transition using LDV. However, the LDV system has a drawback known as the “loading effect,” which inadvertently heats the beam during measurements, as discussed in [25]. We employ digital camera imaging of the beam to assess the bias on the measured $V_{\mathrm{ST}}$  and $V_{\mathrm{SB}}$  values and compare the results obtained with and without LDV. It is important to note that the image processing method for detecting ST and SB events is less precise than the LDV approach. Therefore, we extended the period to 60 seconds in these experiments to achieve more accurate results. This extension was necessary because of the lower temporal resolution of the video camera, which demanded a longer measurement duration for precise observations. Consequently, the bias due to LDV heating could be more prominent than in our experiments’ case of high-frequency measurements due to a significantly longer measurement period.
The figures in Figure 4(a, b) show the results with “LDV on” and “LDV off,” represented by circles and squares.
Figure 4: (a) Critical snap-through voltage ($V_{\mathrm{ST}}$) and (b) critical snap-back voltage ($V_{\mathrm{SB}}$) as a function of the flow velocity $U$. Circle markers indicate measurements with LDV, while square markers indicate measurements without LDV. (c) Sensitivity of the snap-back voltage $V_{\mathrm{SB}}$, measured as a relative voltage change at increasing flow velocity, normalized by the reference value at $U=0$ m/s. The dashed lines represent a quadratic polynomial fit.
The LDV system in the experiment causes an increase in $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$of approximately 4 V and 6 V, respectively. To further quantify the influence of LDV heating, we examine the relative change in SB voltage, denoted $\Delta V_{\mathrm{SB}}=100\times(V_{\mathrm{SB}}-V_{\mathrm{SB}}|{U=0})/V_{\mathrm{SB}}|{U=0}$ in Figure 4(c). Notably, $\Delta V_{\mathrm{SB}}$ increases from 6% to 10% at an airflow velocity of 12 m/s when the LDV is active. We will discuss this bias further in Section 3.1, where we discuss the behavior of the flow velocity sensor due to Joule heating.